# Temporal Entanglement Transitions in the Periodically Driven Ising Chain

### v1

@@ Abstract
As demonstrated in SM [86], exact time evolution under $\mathcal{H}(t)$ shows good agreement with evolution under $\mathcal{H}_{\text{eff}}$ at higher frequencies, with entanglement entropy, Schmidt gap dynamics, and state fidelity exhibiting near-perfect overlap. Temporal entanglement transitions persist in the effective evolution with identical critical exponent $\nu$, confirming that these phenomena represent genuine features of the driven steady state where the EH develops an intrinsic timescale.
Table 1: Conditions for temporal entanglement transitions across different scenarios.
System
Initial state
(Product/Entangled)
Transitions
Equilibrium TFIM Any No
Driven TFIM Product No
Driven TFIM Entangleda Yes
Conserved‐charge
models
Any Yesb
a See SM for detailed discussion on dependence on initial state. b Only when $\rho_{A}$ is forced between disconnected sectors [51, 52].
Discussion and Conclusion.— We have demonstrated that the Floquet driven Ising chain hosts temporal entanglement transitions characterized by dynamical spontaneous symmetry breaking within the EH. These transitions exhibit universal critical behavior with exponent $\nu=1$, identical to equilibrium quantum criticality in the transverse-field Ising universality class, yet emerge from non-equilibrium dynamics that are decoupled from static criticality, as shown by their persistence across different equilibrium phases ($h_{0}=1.6$ to $2.4$, see SM [86]).
This work establishes that: (1) the ES develops non-analyticities for all $\omega$, with periodic reorganizations at higher frequencies; (2) transitions are diagnosed by synchronized Schmidt gap closure, entanglement echo vanishing, and parity flipping, evidencing dynamical spontaneous symmetry breaking; (3) finite-size scaling reveals universal $\nu=1$ critical behavior across frequencies; (4) at high frequencies, Floquet-Magnus theory provides an effective description accurately capturing the transition; (5) as summarized in Table 1, transitions require both driving and initial entanglement, distinguishing them from transitions found for initial product states [53] (more details in the SM [86]).
@@
These findings establish temporal entanglement transitions as fundamental aspects of non-equilibrium Floquet quantum dynamics. Notably, previous work on Floquet symmetry-protected topological system has demonstrated band crossings in entanglement spectrum [94], however the question of robust universality has been an open question. This is what we have addressed in this work, where our temporal entanglement transitions arise universally for any initially entangled state (topological or trivial) and exhibit nonanalytic reorganization (quantum phase transition) of the EH with Ising‐class critical exponents $\nu=1$. The observed behavior suggests driven systems can host rich critical phenomena as diagnosed through their entanglement dynamics. Temporal entanglement transitions define a class of non-equilibrium criticality distinct from Loschmidt echo singularities and magnetization dynamics (see SM [86]). The exact Jordan-Wigner equivalence between the TFIM and Kitaev chain suggests that these transitions should also manifest in driven topological superconductors, opening pathways to entanglement-based probes of Floquet Majorana physics [95, 96]. The requirement for initial entangled states motivates future exploration of how entanglement resources control critical structures in quantum resource theories [97]. Future directions include extensions to higher dimensions, long-range interactions, and potential realization in cold-atoms or trapped ion experiments [98, 99, 100].
Acknowledgments.— K. G. and R. J. acknowledge financial support by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Grants No. 436382789, and No. 493420525, via large equipment grants (GOEGrid). This work is supported by the Office of Advanced Scientific Computing Research, Exploratory Research for Extreme Scale Science (EXPRESS) program, Office of Science of the U.S. Department of Energy under Award Number DE-SC0026216 (A. P.). This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). A. P. thanks the Kavli Institute for Theoretical Physics (KITP) for its hospitality during the program on “Noise-robust Phases of Quantum Matter,” during which part of this work was completed.

### v2

@@ Abstract
As demonstrated in SM [66], exact time evolution under $\mathcal{H}(t)$ shows good agreement with that under $\mathcal{H}_{\text{eff}}$ at higher frequencies, with entanglement entropy, Schmidt gap dynamics, and state fidelity exhibiting near-perfect overlap. TET persist in the effective evolution with identical critical exponent $\nu$, confirming that these phenomena represent genuine features of the driven steady state where the EH develops an intrinsic timescale.
Table 1: Conditions for TET.
System Drive $\mathbb{Z}_{2}$ Initial state $\mathbb{Z}_{2}$ TET
Equilibrium TFIM No
Driven TFIM Yes No (symmetry-broken) No
Driven TFIM Yes Yes Yesa
Conserved-charge models Any Yesb
a Provided $\rho_{A}(t)$ carries weight in both subsystem-parity sectors, with TET occurring when $w_{+}(t_{c}^{(k)})=w_{-}(t_{c}^{(k)})$ while $\rho_{A},P_{A}]$ remains negligible, see SM Sec. III. Many standard product-state preparations are $\mathbb{Z}_{2}$-breaking and thus excluded, while $\mathbb{Z}_{2}$-symmetric (often entangled) initial states are allowed.
b Only when $\rho_{A}$ is dynamically forced between disconnected sectors [52, 53].
Discussion and Conclusion.— We have demonstrated that the Floquet driven Ising chain hosts temporal entanglement transitions (TET) characterized by dynamical spontaneous symmetry breaking within the EH. These transitions exhibit universal critical behavior with exponent $\nu=1$ for critical initial state ($h_{0}/2=1.0$), identical to equilibrium quantum criticality in the transverse-field Ising universality class, yet emerge from non-equilibrium dynamics that are decoupled from static criticality, as shown by their persistence across different equilibrium phases ($h_{0}=1.6$ to $2.4$, see SM [66]).
This work establishes that: (1) the ES develops non-analyticities for all $\omega$, with periodic reorganizations at higher frequencies; (2) transitions are diagnosed by synchronized Schmidt gap closure, entanglement echo vanishing, and parity flipping, evidencing dynamical spontaneous symmetry breaking; (3) finite-size scaling for the critical initial state reveals universal $\nu=1$ critical behavior across frequencies; (4) at high frequencies, Floquet-Magnus theory provides an effective description accurately capturing the transition; (5) TET occur precisely when the sector weights become equal, $w_{+}(t_{c}^{(k)})=w_{-}(t_{c}^{(k)})$, provided the driven Hamiltonian and initial state preserve global $\mathbb{Z}_{2}$ symmetry (Table 1; also see SM, Sec. III [66]). Experimental access to these transitions is feasible: the key observables (Schmidt gap, parity expectations, entanglement echo) require only low-lying eigenstates of the EH, measurable via randomized measurements [75, 76] or Ramsey spectroscopy [77].
@@
These findings establish TET as fundamental aspects of non-equilibrium Floquet quantum dynamics. Notably, previous work on Floquet symmetry-protected topological system has demonstrated band crossings in entanglement spectrum [78], however the question of robust universality has been an open question. This is what we have addressed in this work, where our TET arise whenever the driven Hamiltonian and initial state respect global $\mathbb{Z}_{2}$, with critical times determined by sector-weight crossings $w_{+}(t_{c}^{(k)})=w_{-}(t_{c}^{(k)})$, yielding a nonanalytic reorganization (quantum phase transition) of the EH with Ising-class critical exponents $\nu=1$. TET define a class of non-equilibrium criticality distinct from Loschmidt echo singularities and magnetization dynamics (see SM [66]). The exact Jordan-Wigner equivalence between the TFIM and Kitaev chain suggests that these transitions should also manifest in driven topological superconductors, opening pathways to entanglement-based probes of Floquet Majorana physics [79, 80]. The requirement of symmetry-preserving initial conditions with critical times selected by sector-weight degeneracies motivates future exploration of how symmetry and entanglement resources jointly control non-equilibrium criticality structures in quantum resource theories [81]. Beyond the TFIM, the criterion (global symmetry, symmetric initial state, sector-weight equality $w_{+}(t)=w_{-}(t)$) suggests analogous transitions in other driven systems with discrete symmetries, including higher-spin chains, symmetry-protected topological phases, and lattice gauge theories, with experimental prospects in cold-atom and trapped-ion platforms [82, 83, 84, 75].
Acknowledgments.— K. G. and R. J. acknowledge financial support by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Grants No. 436382789, and No. 493420525, via large equipment grants (GOEGrid). This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research via the Exploratory Research for Extreme Scale Science (EXPRESS) program under Award Number DE-SC0026216. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). A. P. thanks the Kavli Institute for Theoretical Physics (KITP) for its hospitality during the program on “Noise-robust Phases of Quantum Matter”, during which part of this work was completed.
@@
Disclaimer.— This report was prepared as an account of work sponsored by an agency of the United States Government. Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights. Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof.